OpenStudy (anonymous):

Would the answer to |x-2|+4=3 be that there isn't a solution?

OpenStudy (anonymous):

i think the answer would be x=1

OpenStudy (anonymous):

because the equation can still be solved

OpenStudy (anonymous):

|x-2|+4=3 x-2+4=3 or x-2+4=-3 x+2=3 or x+2=-3 x=3-2 or x=-3-2 x=1 or x=-5

OpenStudy (anonymous):

there are no values of x for which the original equation would be true. So no solutions.

OpenStudy (anonymous):

how when you have to subtract 4 from both sides which would give you x=3 and x=-5. Neither of these solutions would work